# Millimeter-wave Radio SLAM: End-to-End Processing Methods and Experimental Validation

### v1

@@ IV-C Regularized Robust Least Squares Estimator
Maximizing the posterior as given in (17), is equivalent to minimizing the negative log-likelihood
@@
We then set the next estimate $\hat{{\mathbf{x}}}^{(j+1)}$ to be equal to the minimum, which gives
@@ IV-D1 UE Initialization Without Prior Information
The challenge in initializing the UE state using the LoS measurement is the unknown clock bias $B_{\textrm{UE}}$. One can consider different trial values of $B_{\textrm{UE}}$ over a range of $[{B}_{\textrm{UE,min}},\,{B}_{\textrm{UE,max}}]$ and for each trial value $\hat{B}_{\textrm{UE}}$, we form an augmented measurement $\check{{\mathbf{z}}}=[\hat{\tau}_{1},\,\hat{\phi}_{1},\,\hat{\theta}_{1},\,\hat{B}_{\textrm{UE}}]^{\top}$ and covariance $\check{{\mathbf{R}}}_{1}=\textrm{blkdiag}({\mathbf{R}}_{1},\sigma_{\hat{B}_{\textrm{UE}}}^{2})$, in which $\sigma_{\hat{B}_{\textrm{UE}}}^{2}$ is variance of the bias which is set higher than variance of the delay estimates. Then, the mean and covariance of the UE prior are given by
@@
in which $\hat{x}_{1}=(\hat{\tau}_{1}c-\hat{B}_{\textrm{UE}})\cos(\alpha_{\textrm{BS}}+\hat{\phi}_{1})$ and $\hat{y}_{1}=(\hat{\tau}_{1}c-\hat{B}_{\textrm{UE}})\sin(\alpha_{\textrm{BS}}+\hat{\phi}_{1})$. After computing the moments using (32), the landmark locations can be computed as presented in Section IV-D2. Then the trial with lowest cost according to (21), which also involves the landmarks, is selected as the UE prior. Due to computational reasons, we solve the described problem using constrained nonlinear optimization [46] for which the problem can be defined as
@@ IV-D2 Landmark Initialization
The affine approximation of the measurement likelihood in (18) around the UE prior reads
@@
where ${\mathbf{J}}_{n}(\boldsymbol{\mu}_{\mathbf{s}})=\nabla_{{\mathbf{s}}}{\mathbf{h}}_{n}({\mathbf{x}})|_{{\mathbf{s}}=\boldsymbol{\mu}_{\mathbf{s}}}$ is the Jacobian of ${\mathbf{h}}_{n}({\mathbf{x}})$ with respect to state ${\mathbf{s}}$ evaluated at $\boldsymbol{\mu}_{\mathbf{s}}$. Hence, the likelihood can be approximated as
@@ V Indoor Environment, Tools and Data
We consider a modern indoor environment at the Hervanta Campus of Tampere University, Finland, located in the so-called Campus Arena building. The environment is illustrated in terms of a floor plan and an actual photograph in Fig. 4, consisting of a fairly large partially open space containing a number of different landmarks such as columns, short walls, booths, and so forth. The BS is located in a narrow annex $3$m wide while the UE moves in the area with a trajectory shown in Fig. 4. Furthermore, $4\times 16$ planar antenna arrays are considered at both the transmitting and receiving ends, with azimuth 3 dB beamwidth of around ${}^{\circ}$. Such antenna system assumption is certainly implementation feasible at BS end while for UEs the current mmWave implementations consider somewhat reduced antenna counts and thus broader beams. The considered PRS bandwidth is 400 MHz following [34].
@@ V-B Measurement Setup and Data
The complete 60 GHz I/Q measurement data set as well as the corresponding processed channel parameter data set, together with supportive scripts are all openly shared along the final published article.
@@ VI-A Angle Extraction Results with Ray Tracing Data
Figure 5: Illustration of angle extraction methods in NLoS (upper row) and LoS (lower row) UE locations. (a) Cancellation method [15] for $N_{\text{peaks}}=10$, $\epsilon=0.05$ (NLoS) and $0.1$ (LoS); (b) CFAR [40] for $N_{\text{TB}}=15$ and $P_{\text{FA}}=0.002$ (NLoS) and $0.12$ (LoS); (c) Proposed SVD method for $p$ of $99.9\%$ (NLoS) and $99.99\%$ (LoS). Extracted path angles are marked with colored dots, side-lobe false detections with arrows, and missed detections with squares.
The method in [15] is parameterised with the so-called support power increment, $\epsilon$, and the maximum number of searched peaks, $N_{\text{peaks}}$. We additionally apply power thresholding to reduce the amount of false detections, and thereon to have as fair comparison as possible. The CFAR technique, in turn, uses training cells of size $N_{\text{TB}}$ around the target cell to estimate the local noise level to reach the given false alarm probability $P_{\text{FA}}$. The Matlab 2D CFAR detector implementation from phased array system toolbox was used, while further clustering was also applied to the results as CFAR method yields easily multiple AoD-AoA estimates per one peak.
@@ VI-A1 Qualitative Comparison
Figure 6: Quantitative assessment of different angle extraction methods. (a) GOSPA metric at different UE/RX locations for cancellation-based method [15] with $\epsilon=0.1$ ( ) and $\epsilon=0.05$ ( ), for CFAR method [40] with $P_{\text{FA}}=0.12$ ( ) and $P_{\text{FA}}=0.002$ ( ), and for the proposed SVD-based method with $p=99.9\%$ ( ) and $p=99.99\%$ ( ). (b) Corresponding average GOSPA metrics for all methods. (c) Average sidelobe detection metric for all methods.
@@ VI-A2 Quantitative Comparison
The quantitative comparison between the three methods in terms of GOSPA and sidelobe detection metric are presented in Fig. 6, covering the whole UE/RX trajectory shown in Fig. 4, while also noting the options for the different parametrizations. In terms of the GOSPA metric, the proposed SVD-based method clearly outperforms the reference methods. Additionally, in terms of the SLFD metric, the proposed method outperforms that in [15] by a large margin. The CFAR with $P_{\text{FA}}=0.002$ is closer in performance, however, having clear challenges to detect the actual multipaths as shown in Fig. 5 already.
In summary, the proposed SVD approach outperforms the prior-art benchmark methods. As demonstrated by the ray-tracing results, it offers robust performance in both LoS and NLoS scenarios, while is also having built-in mechanism to suppress the impacts of the unavoidable antenna sidelobes.
@@ VI-B SLAM Results with Measurement Data
Next, the actual end-to-end SLAM results are provided and analyzed, utilizing the true measurement data. We use the proposed SVD method for angle estimation, and assume $p=99.9\%$. The ToA estimation is carried out as described in (13)–(15), while using brute force search to solve (14)
The proposed SLAM method is benchmarked with respect to two other SLAM approaches [28, 31]. The first benchmark is a geometry based snapshot SLAM algorithm [28] referred to as BM1 and the second is a recursive probability hypothesis density (PHD)-SLAM filter [31] referred to as BM2. In the following, the time delay and clock bias are converted to distance for convenience, and the clock bias is emulated to evolve according to a random walk model with variance $\sigma_{B}^{2}=1\,\textrm{m}^{2}$. The measurement noise covariance used in the experiments is ${\mathbf{R}}_{n}=\textrm{diag}([0.3\,\textrm{m},\;3\,\textrm{deg},\;3\,\textrm{deg}]^{2}),\,\forall\,n$. The mean and covariance of the UE prior for the proposed method are $\boldsymbol{\mu}_{\mathbf{s}}=\boldsymbol{\mu}_{{\mathbf{s}},\textrm{prev}}$ and $\boldsymbol{\Sigma}_{{\mathbf{s}}{\mathbf{s}}}=\mathbf{I}_{4}$, in which $\boldsymbol{\mu}_{{\mathbf{s}},\textrm{prev}}$ denotes the UE state estimate at the previous measurement position. In the first measurement position, which is located at $[0.55,\,-2.75]^{\top}$, there is no prior and estimation is possible because the LoS signal exists. The PHD-SLAM filter is implemented using $1000$ particles, the UE state is modeled to evolve according to a random walk model and the process noise is set to ${\mathbf{Q}}=\mathbf{I}_{4}$ which is the same as covariance of the UE prior. The PHD-SLAM filter is initialized using the proposed snapshot SLAM algorithm. Since BM1 is sensitive to outliers, measurements labeled as outliers  [^1] Under the Gaussian assumption, the quadratic error in (22) follows a $\chi^{2}$ distribution with three degrees of freedom and if $q_{n}({\mathbf{x}})>T_{h}$, the measurement is labeled an outlier. The quadratic error is evaluated using the ground truth UE state and $T_{h}\approx 16.3$ is computed by choosing tail probability $0.001$ followed by evaluating the inverse cumulative distribution of the $\chi^{2}$ distribution at $0.999$. are removed from the data for BM1 unless otherwise stated.
@@ VI-B1 Qualitative Comparison
The example mapping and localization performance of the algorithms are visualized in Fig. 7. Qualitative analysis indicates that the map and UE position estimates are more accurate with BM2 and the proposed SLAM algorithm than with BM1. For BM1, the estimate is close to the ground truth and the estimated landmark locations are inline with the map in many measurement positions. In several locations however, the estimates are very inaccurate and there are two primary reasons for this. First of all, the method requires four NLoS signals so that the system is identifiable and this criterion is not satisfied at every measurement position. For example, when the UE is located at ${\mathbf{p}}_{\textrm{UE}}=[-5.5,\,-2.3]^{\top}$, there are only three propagation paths meaning that the system is underdetermined resulting in an inaccurate estimates as illustrated in Fig. 7. The second reason is that in several measurement positions, the cost function of BM1 is multimodal and the global minima is not the one closest to the ground truth. The proposed method and BM2 can operate in mixed LoS/NLoS conditions and the prior or posterior from the previous time step can be viewed as a regularization term which constrains the posterior update so that the system state is identifiable at every measurement position. Moreover, the resulting estimates for BM2 and the proposed method can be viewed as a weighted average of the evidence provided by the data and the regularization term. Thus, the estimate is expected to be close to the ground truth as long as the prior is not biased and the covariance captures the underlying uncertainties correctly. Lastly, both BM2 and the proposed approach result in sufficient SLAM performance despite the measurement data being corrupted by outliers as shown in Fig. 7.
@@ VI-B2 Quantitative Comparison
TABLE I: Performance summary of the different estimators
@@
BM1 $1.70\pm 1.36$ $29.17\pm 28.98$ $1.85\pm 1.86$ $13.43$
$*{}^{*}$ $1.51\pm 1.23$ $14.26\pm 14.16$ $0.92\pm 0.86$ $1.39$
BM2 $0.55\pm 0.23$ $2.56\pm 2.45$ $0.46\pm 0.40$ $141.12$
Proposed $0.55\pm 0.32$ $2.26\pm 2.18$ $0.52\pm 0.50$ $1.60$
*The PHD-SLAM filter is implemented with $10$ particles.
The algorithms are next evaluated quantitatively while since the ground truth landmark locations are unknown, the mapping accuracy is excluded which a common practice in SLAM when using experimental data. The performance metrics are tabulated in Table I in which the position, heading and bias errors are computed using the RMSE and standard deviation (STD). Even without outliers, BM1 results in unsatisfactory performance due to the reason discussed above. The other two methods outperform BM1 and have comparative performance among each other. However, the computational complexity of BM2 is two orders of magnitude higher than with the proposed method. One could decrease the computational complexity by using less particles (see $*{}^{*}$ in Table I) but at the same time, the accuracy of the filter degrades notably. Thus, a major advantage of the proposed method is that it combines high accuracy together with low computational overhead.
@@
Next, we evaluate the different algorithms with and without outliers, and in LoS and NLoS conditions. The results are summarized in Fig. 8 and in general, the accuracy of all methods is better when the LoS is available. The performance of BM1 degrades significantly if the data contains outliers, whereas the performance of BM2 and the proposed method only degrade slightly. This indicates that both methods are capable of handling noisy measurements that are not inline with the models. The mechanisms how the two methods deal with outliers are quite different. The PHD-SLAM filter associates the outliers to clutter so that they do not affect the UE estimate, whereas the proposed method inflates the covariance according to (29) such that outliers are given less trust thus diminishing their impact. Interestingly, the proposed method outperforms BM2 in LoS conditions and vice versa in NLoS conditions. For the proposed method and when the LoS signal exists, the hypothesis that minimizes the cost function is typically the one for which the prior is computed using the LoS and the estimate is computed only from the evidence provided by the data. In this particular scenario, in which the prior uncertainty and the process noise of the dynamic model used by BM2 are high, relying solely on the measurements is beneficial. In NLoS conditions and when the proposed method cannot be initialized using the measurements, filtering is beneficial and BM2 outperforms the proposed method – however, at the expense of substantially higher complexity.

### v2

@@ IV-C Regularized Robust Least Squares Estimator
Maximizing the posterior as given in (19), is equivalent to
@@
In order to minimize $L({\mathbf{x}})$, we then set the next estimate $\hat{{\mathbf{x}}}^{(j+1)}$ to be equal to the minimum, which gives
@@ IV-D1 UE Initialization Without Prior Information
The challenge in initializing the UE state using the LoS measurement is the unknown clock bias $B_{\textrm{UE}}$. One can consider different trial values of $B_{\textrm{UE}}$ over a range of $[{B}_{\textrm{UE,min}},\,{B}_{\textrm{UE,max}}]$ [^1] We can define the following inequality, $d_{\textrm{min}}\leq c\hat{\tau}^{b}_{1}-B_{UE}\leq d_{\textrm{max}}$, in which $d_{\textrm{min}}$ and $d_{\textrm{max}}$ define the minimum and maximum propagation distances of the LoS path. From the inequality, the bias range can be computed as $B_{\textrm{UE,min}}=c\hat{\tau}^{b}_{1}-d_{\textrm{max}}$ and $B_{\textrm{UE,max}}=c\hat{\tau}^{b}_{1}-d_{\textrm{min}}$. and for each trial value $\hat{B}_{\textrm{UE}}$, we form an augmented measurement $\check{{\mathbf{z}}}=[\hat{\tau}^{b}_{1},\,\hat{\phi}_{1},\,\hat{\theta}_{1},\,\hat{B}_{\textrm{UE}}]^{\top}$ and covariance $\check{{\mathbf{R}}}_{1}=\textrm{blkdiag}({\mathbf{R}}_{1},\sigma_{\hat{B}_{\textrm{UE}}}^{2})$, in which $\sigma_{\hat{B}_{\textrm{UE}}}^{2}$ is variance of the bias which is set higher than variance of the delay estimates. Then, the mean and covariance of the UE prior are given by
@@
in which $\hat{x}_{1}=(c\hat{\tau}^{b}_{1}-\hat{B}_{\textrm{UE}})\cos(\alpha_{\textrm{BS}}+\hat{\phi}_{1})$ and $\hat{y}_{1}=(c\hat{\tau}^{b}_{1}-\hat{B}_{\textrm{UE}})\sin(\alpha_{\textrm{BS}}+\hat{\phi}_{1})$. After computing the moments using (34), the landmark locations can be computed as presented in Section IV-D2. Then the trial with lowest cost according to (23), which also involves the landmarks, is selected as the UE prior. Due to computational reasons, we solve the described problem using constrained nonlinear optimization [48] for which the problem can be defined as
@@ IV-D2 Landmark Initialization
The first order Taylor series based linear approximation of the measurement likelihood in (20) around the UE prior reads
@@
where ${\mathbf{h}}_{n}({\mathbf{x}})$ is evaluated at ${\mathbf{x}}=[\boldsymbol{\mu}_{\mathbf{s}}^{\top},{\mathbf{m}}_{n}^{\top}]^{\top}$, and ${\mathbf{J}}_{n}(\boldsymbol{\mu}_{\mathbf{s}})=\nabla_{{\mathbf{s}}}{\mathbf{h}}_{n}({\mathbf{x}})|_{{\mathbf{s}}=\boldsymbol{\mu}_{\mathbf{s}}}$ is the Jacobian of ${\mathbf{h}}_{n}({\mathbf{x}})$ with respect to state ${\mathbf{s}}$ computed at $\boldsymbol{\mu}_{\mathbf{s}}$. Following the derivations in [49, Eqs. (5.6) - (5.13)], the likelihood can be approximated as
@@ V Indoor Environment, Tools and Data
We consider a modern indoor environment at the Hervanta Campus of Tampere University, Finland, located in the so-called Campus Arena building. The environment is illustrated in terms of a floor plan and an actual photograph in Fig. 4, consisting of a fairly large partially open space containing a number of different landmarks such as columns, short walls, booths, and so forth. The BS is located in a narrow annex $3$m wide while the UE moves in the area with a trajectory shown in Fig. 4. Furthermore, $4\times 16$ planar antenna arrays are considered at both the transmitting and receiving ends, with azimuth 3 dB beamwidth of around 10. Such antenna system assumption is certainly implementation feasible at BS end while for UEs the current mmWave implementations consider somewhat reduced antenna counts and thus broader beams. The considered PRS bandwidth is 400 MHz following [36].
@@ V-B Measurement Setup and Data
The complete 60 GHz I/Q measurement data set as well as the corresponding processed channel parameter data set, together with supportive scripts are all openly shared along this article – refer to the first page footnote for further details.
@@ VI-A Angle Extraction Results with Ray Tracing Data
Figure 5: Illustration of angle extraction methods in NLoS (upper row) and LoS (lower row) UE locations. (a) Cancellation method [15] for $N_{\text{peaks}}=10$, $\epsilon=0.05$ (NLoS) and $0.1$ (LoS); (b) CFAR [42] for $N_{\text{TB}}=15$ and $P_{\text{FA}}=0.002$ (NLoS) and $0.12$ (LoS); (c) Proposed SVD method for $p$ of $99.9\%$ (NLoS) and $99.99\%$ (LoS), and with $\beta_{\text{th}}$ $10\%$ above the noise floor. Extracted path angles are marked with colored dots, side-lobe false detections with arrows, and missed detections with squares.
The method in [15] is parameterised with the so-called support power increment, $\epsilon$, and the maximum number of searched peaks, $N_{\text{peaks}}$. We additionally apply power thresholding to reduce the amount of false detections, and thereon to have as fair comparison as possible. The CFAR technique, in turn, uses training cells of size $N_{\text{TB}}$ around the target cell to estimate the local noise level to reach the given false alarm probability $P_{\text{FA}}$. The Matlab 2D CFAR detector implementation from phased array system toolbox was used, while further clustering was also applied to the results as CFAR method yields easily multiple AoD-AoA estimates per one peak. Finally, the proposed SVD-based method is parametrized as follows. For the power ratio parameter $p$, we consider the values of $p=99\%,99.9\%$ and $99.99\%$. The power thresholding related parameter $\beta_{\text{th}}$ is, in turn, set at $10\%$ above the prevailing noise floor. This value is designated with notation $\beta_{\text{th}}^{*}$ while additional complementary results where the value of $\beta_{\text{th}}$ is varied are also provided.
@@ VI-A1 Qualitative Comparison
Figure 6: Quantitative assessment of different angle extraction methods. (a) GOSPA metric at different UE/RX locations for cancellation-based method [15] with $\epsilon=0.1$ ( ) and $\epsilon=0.05$ ( ), for CFAR method [42] with $P_{\text{FA}}=0.12$ ( ) and $P_{\text{FA}}=0.002$ ( ), and for the proposed SVD-based method with $p=99\%$,( ) $p=99.9\%$ ( ), and $p=99.99\%$ ( ), and with $\beta_{\text{th}}$ $10\%$ above the noise floor. (b) Corresponding average GOSPA metrics for all methods. (c) Average sidelobe detection metric for all methods.
@@ VI-A2 Quantitative Comparison
The quantitative comparison between the three methods in terms of GOSPA and sidelobe detection metric are presented in Fig. 6, covering the whole UE/RX trajectory shown in Fig. 4, while also noting the options for the different parametrizations. In terms of the GOSPA metric, the proposed SVD-based method clearly outperforms the reference methods. Additionally, in terms of the SLFD metric, the proposed method outperforms that in [15] by a large margin. The CFAR with $P_{\text{FA}}=0.002$ is closer in performance, however, having clear challenges to detect the actual multipaths as shown in Fig. 5 already. It can also be observed that for the proposed method, the GOSPA metric decreases with the increase of $p$ due to the increased number of recovered multipath components. At the same time, the sidelobe detection rate slightly increases with the number of detected components, however, the sidelobe detection metric stays generally low for all considered values of $p$ compared to the reference methods.
We next provide further complementary assessment of the SVD-based method in terms of the impact of the power thresholding parameter $\beta_{\text{th}}$. The results are shown in Fig. 7), with $\beta_{\text{th}}^{*}$ denoting the baseline value used otherwise in the article. We can observe that the exact value of the power thresholding is impacting the performance only when working under a vary large value of the power ratio parameter $p$ (particularly $p=99.99\%$). This is because in such case, the baseline SVD processing passes through a large number of components, some of which being most likely noise already. Overall, the results in Fig. 7 show the robust behavior and performance of the proposed scheme.
Figure 7: GOSPA metric for the proposed SVD-based method for $p=99\%$ ( ), $p=99.9\%$ ( ), and $p=99.99\%$ ( ) with different values for the power thresholding parameter $\beta_{\text{th}}$ relative to the baseline value $\beta_{\text{th}}^{*}$.
In summary, we conclude that the proposed SVD approach outperforms the prior-art benchmark methods. As demonstrated by the ray-tracing results, it offers robust performance in both LoS and NLoS scenarios, while is also having built-in mechanism to suppress the impacts of the unavoidable antenna sidelobes.
@@ VI-B SLAM Results with Measurement Data
(a) BM1 [28]
(b) BM2 [31]
(c) Proposed SLAM approach
Figure 9: Visual illustration of the SLAM performance in the reference case where the UE and BS are mutually synchronized. For other details, refer to the caption of Fig. 8).
Next, the actual end-to-end SLAM results are provided and analyzed, utilizing the true measurement data. We use the proposed SVD method for angle estimation, and assume $p=99\%$. The ToA estimation is carried out as described in (14)–(17), while using brute force search to solve (16).
The proposed SLAM method is benchmarked with respect to two other SLAM approaches [28, 31]. The first benchmark is a geometry based snapshot SLAM algorithm [28] referred to as BM1 and the second is a recursive probability hypothesis density (PHD)-SLAM filter [31] referred to as BM2. In the following, the time delay and clock bias are converted to distance for convenience, and the clock bias is emulated to evolve according to a random walk model with variance $\sigma_{B}^{2}=1\,\textrm{m}^{2}$. The measurement noise covariance used in the experiments is ${\mathbf{R}}_{n}=\textrm{diag}([0.3\,\textrm{m},\;3\,\textrm{deg},\;3\,\textrm{deg}]^{2}),\,\forall\,n$. Recall that no prior knowledge of the map is assumed as discussed in Section IV-D. The mean and covariance of the UE prior for the proposed method are $\boldsymbol{\mu}_{\mathbf{s}}=\boldsymbol{\mu}_{{\mathbf{s}},\textrm{prev}}$ and $\boldsymbol{\Sigma}_{{\mathbf{s}}{\mathbf{s}}}=\mathbf{I}_{4}$, in which $\boldsymbol{\mu}_{{\mathbf{s}},\textrm{prev}}$ denotes the UE state estimate at the previous measurement position. In the first measurement position, which is located at $[0.55,\,-2.75]^{\top}$, there is no prior and estimation is possible because the LoS signal exists. The UE initialization algorithm presented in Section IV-D1 uses $d_{\textrm{min}}=1\text{ m}$ and $d_{\textrm{max}}=20\text{ m}$ in the constrained nonlinear optimization problem defined in Equation (35). The benchmark PHD-SLAM filter considers three types of landmarks [14]: i) BS for the LoS path; ii) virtual anchors (VAs) for large reflecting surfaces; and iii) scattering points (SPs) for small objects. However, the VAs are converted to SPs using [14, Eq. (42)] in the following illustrations so that the map of the PHD-SLAM filter is directly comparable to the maps of the snapshot SLAM algorithms. The benchmark PHD-SLAM filter uses similar parameter values as in [31, 54], while the values were slightly tuned to maximize the performance in the considered experiment. The PHD-SLAM filter is implemented using $1000$ particles, the UE state is modeled to evolve according to a random walk model and the process noise is set to ${\mathbf{Q}}=\mathbf{I}_{4}$ which is the same as covariance of the UE prior. The PHD-SLAM filter is initialized using the proposed snapshot SLAM algorithm. Since BM1 is sensitive to outliers, measurements labeled as outliers [^2] Under the Gaussian assumption, the quadratic error in (24) follows a $\chi^{2}$ distribution with three degrees of freedom and if $q_{n}({\mathbf{x}})>T_{h}$, the measurement is labeled an outlier. The quadratic error is evaluated using the ground truth UE state and $T_{h}\approx 16.3$ is computed by choosing tail probability $0.001$ followed by evaluating the inverse cumulative distribution of the $\chi^{2}$ distribution at $0.999$. are removed from the data for BM1 unless otherwise stated.
@@ VI-B1 Qualitative Comparison
The example mapping and localization performance of the algorithms are visualized in Fig. 8. Qualitative analysis indicates that the map and UE position estimates are more accurate with BM2 and the proposed SLAM algorithm than with BM1. For BM1, the estimate is close to the ground truth and the estimated landmark locations are inline with the map in many measurement positions. In several locations however, the estimates are very inaccurate and there are two primary reasons for this. First of all, the method requires four NLoS signals so that the system is identifiable and this criterion is not satisfied at every measurement position. For example, when the UE is located at ${\mathbf{p}}_{\textrm{UE}}=[4.06,\,-5.70]^{\top}$, there are only three propagation paths meaning that the system is underdetermined resulting in an inaccurate estimates as illustrated in Fig. 8. The second reason is that in several measurement positions, the cost function of BM1 is multimodal and the global minima is not the one closest to the ground truth. The proposed method and BM2 can operate in mixed LoS/NLoS conditions and the prior or posterior from the previous time step can be viewed as a regularization term which constrains the posterior update so that the system state is identifiable at every measurement position. Moreover, the resulting estimates for BM2 and the proposed method can be viewed as a weighted average of the evidence provided by the data and the regularization term. Thus, the estimate is expected to be close to the ground truth as long as the prior is not biased and the covariance captures the underlying uncertainties correctly. Lastly, both BM2 and the proposed approach result in sufficient SLAM performance despite the measurement data being corrupted by outliers as shown in Fig. 8.
Since both the TX and RX were controlled by the same PC, with an opportunity to utilize also a dedicated synchronization cable allowing to synchronize to a common clock, we can also analyze the reference performance of the algorithms with known clock bias $B_{\textrm{UE}}$. Example performance is illustrated in Fig. 9 while the results are further elaborated and discussed in Section VI-B3.
@@ VI-B2 Quantitative Comparison
TABLE I: Performance summary of the different estimators with and without clock bias
@@
BM1 $2.83\pm 2.24$ $32.05\pm 32.13$ $2.99\pm 3.03$ $41.44$
BM2 $0.56\pm 0.26$ $2.85\pm 2.79$ $0.55\pm 0.37$ $77.41$
BM2 $2.86\pm 2.37$ $35.19\pm 34.92$ $1.87\pm 1.83$ $1.16$
Proposed $0.56\pm 0.33$ $2.30\pm 2.26$ $0.54\pm 0.51$ $1.44$
BM1 $0.34\pm 0.16$ $2.32\pm 2.33$- $0.37$
BM2 $0.26\pm 0.14$ $2.24\pm 2.21$- $73.83$
Proposed $0.32\pm 0.16$ $1.87\pm 1.79$- $0.15$
BM2 implemented with $10$ particles. Known clock bias $B_{\textrm{UE}}$.
The algorithms are next evaluated quantitatively while since the ground truth landmark locations are unknown, the mapping accuracy is excluded which a common practice in SLAM when using experimental data. The performance metrics are tabulated in Table I in which the position, heading and bias errors are computed using the RMSE and standard deviation (STD). Even without outliers, BM1 results in unsatisfactory performance due to the reason discussed above. The other two methods outperform BM1 and have comparative performance among each other. However, the computational complexity of BM2 is two orders of magnitude higher than with the proposed method. One could decrease the computational complexity by using less particles (see BM2 in Table I) but at the same time, the accuracy of the filter degrades notably. Thus, a major advantage of the proposed method is that it combines high accuracy together with low computational overhead.
@@
Next, we evaluate the different algorithms with and without outliers, and in LoS and NLoS conditions. The results are summarized in Fig. 10. The performance of BM1 degrades significantly if the data contains outliers, whereas the performance of BM2 and the proposed method only degrade slightly. This indicates that both methods are capable of handling noisy measurements that are not inline with the models. The mechanisms how the two methods deal with outliers are quite different. The PHD-SLAM filter associates the outliers to clutter so that they do not affect the UE estimate, whereas the proposed method inflates the covariance according to (31) such that outliers are given less trust thus diminishing their impact. Interestingly, the proposed method outperforms BM2 in LoS conditions and vice versa in NLoS conditions. For the proposed method and when the LoS signal exists, the hypothesis that minimizes the cost function is typically the one for which the prior is computed using the LoS and the estimate is computed only from the evidence provided by the data. In this particular scenario, in which the prior uncertainty and the process noise of the dynamic model used by BM2 are high, relying solely on the measurements is beneficial. In NLoS conditions and when the proposed method cannot be initialized using the measurements, filtering is beneficial and BM2 outperforms the proposed method – however, at the expense of substantially higher complexity.
@@
VI-B3 SLAM Performance with Known Clock Bias
The estimated angles determine the problem geometry up to a scaling which is defined by the unknown clock bias. Thus, the SLAM problem is notably easier to solve if the clock bias is known. The experimental arrangements allowed for an additional dedicated synchronization cable between the TX and RX entities, such that we could also evaluate the reference system performance with perfect synchronization. Example performance with known clock bias was already illustrated in Fig. 9 while the corresponding quantitative performance metrics are tabulated in Table I. As can be observed, the performance improves significantly with all methods when comparing to the practical case with unknown clock bias. Perfect synchronization improves the performance of BM1 the most. Since $B_{\textrm{UE}}$ is known, only three propagation paths are required to solve the problem and this criterion is satisfied at every measurement position. The cost function is also unimodal and the minima are close to the corresponding ground truth in every measurement position. In addition, the computational complexity of the algorithm is much lower since the problem is solved by finding the minimum cost over $\alpha_{\textrm{UE}}$, whereas with unknown clock bias, the minimum cost is found by jointly optimizing $\alpha_{\textrm{UE}}$ and $B_{\textrm{UE}}$. With unknown clock bias, the underlying likelihood function that is used to update the particle weights of BM2 is commonly multimodal and the filter can converge to the wrong solution. On the other hand, the likelihood function is generally unimodal if $B_{\textrm{UE}}$ is known so that the filter more frequently converges to the correct solution. The computational complexity of BM2 only reduces slightly since the only notable difference is that the UE state dimension reduces from four to three. Also with the proposed algorithm the performance improves significantly and the underlying cause is the same as with BM1 and BM2, that is, the likelihood function is typically unimodal resulting in an objective function that has a minimum close to the ground truth. Furthermore, the initialization of the proposed algorithm simplifies if the TX and RX entities are synchronized since mean and covariance of the UE prior are directly given by (34) and optimization over different trial values of $\hat{B}_{\textrm{UE}}$ is not required. Thanks to this straightforward initialization procedure and faster convergence of the Gauss-Newton algorithm, the computational complexity of the algorithm is reduced by an order of magnitude.
Figure 11: The numbers of inliers and outliers in each measurement position and with different power ratio $p$ values. Minimum number of inliers required by BM1 to solve SLAM is illustrated with the black dashed line.
VI-B4 Impact of Power Ratio to SLAM Accuracy
Finally, we evaluate and show the SLAM performance using channel estimates that are obtained with three different values for the power ratio parameter $p$ in the proposed SVD method. As already discussed in Section VI-A2, increasing $p$ grows the rank of the SVD-based approximation and as an outcome, the number of estimated paths increases. However, the number of resolvable propagation paths cannot be grown indefinitely since a limited number of physical propagation paths exist and power threshold $\beta_{\text{th}}$ truncates channel estimates for which the power is close to the noise floor. Higher $p$ values also increase the number of measurements that do not follow the model in (21), such as multi-bounce paths. Such measurements must be treated as unwanted outliers that cannot be utilized by the considered SLAM algorithms. The number of inliers and outliers in each of the $45$ measurement positions is illustrated in Fig. 11. As shown, increasing the power ratio from $99\%$ to $99.9\%$ increases both the number of inliers and outliers. Increasing $p$ even more, the channel estimates of $p=99.9\%$ and $p=99.99\%$ only differ slightly in two measurement positions since already with $p=99.9\%$, the SVD-based channel estimator finds all resolvable propagation paths. The power ratio value has the biggest impact on BM1. Recall that four or more propagation paths are required to solve the SLAM using BM1 and if $p=99\%$, this criterion is not satisfied at measurement position number $14$ and $35$ as illustrated in Fig. 11. Now increasing $p$ also increases the number of inliers such that the system is identifiable at every measurement position which improves the accuracy as tabulated in Table II. The system is always identifiable with the proposed and BM2 methods since the prior constrains the posterior update and the two methods yield similar performance with the tested power ratio values. In general, it is expected that SLAM performance improves as the number of propagation paths increases [27]. At the same time, however, the number of outliers increases which can have a negative impact on SLAM performance and therefore, the accuracy with BM2 and the proposed approach only improves very little (BM2) or not at all.
TABLE II: Position RMSE in meters with different power ratio $p$ values
Estimator $p=99\%$ $p=99.9\%$ $p=99.99\%$
BM1 $2.83$ $1.43$ $1.43$
BM2 $0.56$ $0.55$ $0.54$
Proposed $0.56$ $0.56$ $0.56$
VI-C Limitations and Future Directions
One of the limitations of the proposed BRSRP-based AoA/AoD extraction algorithm is its inability to decouple two multipath components that have the same AoA or AoD, which may lead to a missed detection. This type of misdetections, however, do not largely impact the actual SLAM performance, as in majority of cases such undetected components correspond to multi-bounce paths due to the involved geometry. Nevertheless, one future work direction is extending the SVD-based AoA/AoD extraction to support multi-peak detection and selection. One potential way to pursue such can be performing the peak search directly in the singular vector domain, aided by further assumptions about the shape of the beam responses and ToA-based sidelobe rejection. Alternatively, the overall problem may be addressed through a different technical approach – for example, as a pattern recognition problem solved either through classical methods or potentially through machine learning. On the SLAM and related concepts side, extending the methods to extract information also about the material characteristics of the environment interaction points, on top of their locations, is an interesting future research avenue.